Warren has 40 coins (all nickels, dimes and quarters) worth $4.05. He has 7 more nickels than dimes. How many quarters does Warren have? Could you please show equation?
Thank you

Answer by stanbon(75887)

Warren has 40 coins (all nickels, dimes and quarters) worth $4.05. He has 7 more nickels than dimes. How many quarters does Warren have? Could you please show equation?
-------------------
Equations:
n + d + q = 40 coins
5n+10d+25q = 405 cents
n = d + 7
-------------------
Rearrange and solve the system.
Cheers,
Stan H.